OpenStudy (sonyalee77):

What is the equation of the Line that passes through (4,2) and is parallel to 3x-2y=-6?

OpenStudy (anonymous):

3x-2y=-6 can be written as y=(3x/2)+3

OpenStudy (sonyalee77):

I thought that it was-3/2+8

OpenStudy (anonymous):

Nope. Parallel lines have the same slope. The slope for y=(3x/2)+3 is (3/2).

OpenStudy (sonyalee77):

But none of my multiple choice answers don't. Match

OpenStudy (anonymous):

That's the not equation for it to go through point, (4,2), that's the equation for 3x-2y=-6

OpenStudy (sonyalee77):

A.-3/2x+8 B.3/2x-4 C.1/2x D.2/3x-4

OpenStudy (anonymous):

You have to input the point into another equation. (y-y1)=m(x-x1), where y1 is the point for y, and x1 is the point for x for the point you want. In this case, the point is (4,2)

OpenStudy (anonymous):

y-2=(3/2)(x-4)

OpenStudy (anonymous):

y-2=(3/2)x-6 +2 +2 y=(3/2)x-4

OpenStudy (anonymous):

(3/2) is the slope (m), which is the same for parallel lines.

OpenStudy (anonymous):

the -6 is (3/2) times -4. We distributed it to everything in (x-4).